A machine costing $209,600 with a four-year life and an estimated $16,000 salvage value is installed in Luther Company’s factory on January 1. The factory manager estimates the machine will produce 484,000 units of product during its life. It actually produces the following units: 122,400 in 1st year, 122,400 in 2nd year, 121,000 in 3rd year, 128,200 in 4th year. The total number of units produced by the end of year 4 exceeds the original estimate—this difference was not predicted. (The machine must not be depreciated below its estimated salvage value.)


Required:

Compute depreciation for each year (and total depreciation of all years combined) for the machine under each depreciation method. (Round your per unit depreciation to 2 decimal places. Round your answers to the nearest whole dollar.)

Complete this question by entering your answers in the tabs below.

  • Straight Line
  • Units of Production
  • DDB

Compute depreciation for each year (and total depreciation of all years combined) for the machine under Straight-line depreciation.

Straight-Line Depreciation
Year Depreciation Expense
1
2
3
4
Total
Answer #1
Straight Line Depreciation:
Cost of machine $209,600
Estimated Salvage Value $16,000
Useful life 4
Annual Depreciation=(Cost of machine-Estimated Salvage Value)/(Useful Life)
Annual Depreciation= $48,400 (209600-1600)/4
Year Depreciation Expense
1 $48,400
2 $48,400
3 $48,400
4 $48,400
Total $193,600
UNITS OF PRODUCTION
Cost of machine $209,600
Estimated Salvage Value $16,000
A Depreciable Value =209600-16000 $193,600
B Estimated Units of production 484000
C=A/B Depreciation rate per unit of production $0.40
D E=D*0.4 F=Max. (193600)
Year Production Depreciation Accumulated Depreciation
1 122400 $48,960 48960
2 122400 $48,960 97920
3 121000 $48,400 146320
4 128200 $47,280 193600
Year 4 depreciation calculated such that Accumulated Depreciation =193600
Year Depreciation Expense
1 $48,960
2 $48,960
3 $48,400
4 $47,280
Total $193,600
DDB (DOUBLE DECLINING BALANCE DEPRECIATION)
Depreciation Rate =2*(1/useful Life)=2/4 50%
Depreciation during the year =50%* Book Value at the beginning of Year
A B=A*50% C=A-B
Year Book Value at beginning Depreciation amount Ending Book ValueMin.(16000)
1 $209,600 $104,800 $104,800
2 $104,800 $52,400 $52,400
3 $52,400 $26,200 $26,200
4 $26,200 $10,200 $16,000
Year 4 depreciation is calculated based on year end book value =Salvage Value
Year Depreciation Expense
1 $104,800
2 $52,400
3 $26,200
4 $10,200
Total $193,600